- Describe the process of planning, managing, and controlling quality in projects.

Class - A project has the following key characteristics: (a) start and end date, and (b) delivers a unique product, service, or result. For this final exam, please select a business-type project that you would like to launch and respond to the following items:


• Provide a description of the project.
• How will you convince your manager that this project is worthwhile?
• Does your project have a favorable return on investment (ROI)? Please discuss.
• How will you secure the necessary resources for this project? Please note resources include people, capital, equipment, technology, supplies, physical location, and so on.
• What are the risk factors that you will encounter, and how can you overcome them?
• What is the change management process you will follow for this project? Please make sure to discuss the formal steps.
• What is the process you will use to collect lessons learned? Where will you archive them?
• How will you define success for your project?
Answer #1

The project that I have chosen is the setup of a modern campaign management project. We are moving away from a traditional batch mode campaign system to a real time big data enabled campaign management system which will enable contextual campaign management.

I will convince the manager that the project is worthwhile by creating the ROI , Payback and IRR calculations and showing that payback is short and ROI and IRR are positive and significant. I will also showcase how this project will positively impact the lives of our customers by improving their experience.

The ROI is favorable as the expected return on investment is quite high. The expected return is positive as the expectation is to earn an additional yearly $200 million against an investment of $10 million. giving a ROI of 19 [ Total Profit/Total Investment = (200-10)/10 = 190/10 = 19]

Once the project charter is signed. Detailed project planning begins where plans are made for each of the project management knowledge areas:

1) Procurement Management Plan

2) Risk Management Plan

3) Integration Management Plan

4) Quality Management Plan

5) Time Management Plan

6) Cost Management Plan

7) Scope Management Plan

8) Human Resource Management Plan

9) Communications Management Plan

Planning is done to the most granular level where Work Breakdown Structure (WBS) and WBS dictionary is created, activity list and network diagram is made. Resources are estimated along with time and cost. Basis this critical path is determined and schedule and budget are developed. Then the quality standards, metrics are set and quality processes are determined. Similarly all the other plans are created.

Once all the resources estimated these are procured or hired as per the created plan.